OpenStudy (anonymous):

A kite 50ft above the ground moves horizontally at a speed of 8ft/s. At what rate is the angle between the string and the horizontal decreasing when 300ft of string has been let out?

OpenStudy (anonymous):

d = r •t t = d / r r = d / t Hope this helps, because these are the formulas.

OpenStudy (anonymous):

still dont get how to set up the equation

OpenStudy (anonymous):

300ft is the distance. and 8ft/s is the time, right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

So, use r = d / t.

OpenStudy (anonymous):

300 / 8 = 37.5

OpenStudy (p0sitr0n):

derivate  \[\cos \alpha = \sqrt{a^2-b^2} / a \]